If p/a + q/b + r/c = 0 and a/p + b/q +c/r = 1 then the value of (p/a)2 + (q/b)+ (r/c)= ?

Grade:12

2 Answers

Vijay Mukati
7 years ago
Dear Student,

Using (a+b+c)2 = a2 + b2+ c2 + 2(ab+bc+ca).
final answer will be -2(pqr/abc).

Thanks.
manmath
7 years ago
dear friends
r u want the answer 4r this questoin then see my answer
+ 2(ab+bc+ca).final answer will be -2(pqr/abc).2+ c2 + b2 = a2Using (a+b+c)